# SPICE[Uncaptioned image]: An Automated SWE-Bench Labeling Pipeline for Issue Clarity, Test Coverage, and Effort Estimation

### v4

@@ Abstract
Index Terms:
Data labeling, code LLM, pretraining, finetuning, SWE-bench, benchmark
@@ I Introduction
In the era of Foundation Models (FMs) like Large Language Models (LLMs), high-quality software engineering (SE) datasets fulfill two key roles: (i) they benchmark model performance on real-world SE tasks and (ii) serve as valuable corpora for pretraining and fine-tuning FMs.
@@
A prominent example of a high-quality SE dataset is SWE-bench (SoftWare Engineering benchmark) Verified [1] (SWE-V). SWE-V was designed to assess an FM’s ability to produce a valid code patch that fixes an issue report from a given software repository. More specifically, each instance (row) in SWE-V represents a code repair task and contains three key fields: the issue description (title and body), a reference (golden) solution patch, and corresponding test cases used for evaluation. We refer to similar SE datasets containing these key fields as SWE-bench-like datasets. Notable SWE-bench-like datasets include SWE-Gym [2], Multi-SWE-bench [3], and SWE-bench Multimodal [4]. The tasks in SWE-V span 12 real-world Python projects, including widely used ones such as Django.
@@
In terms of role (i), SWE-V has been extensively used to benchmark FMs as well as agentic scaffolds built on top of those. A leaderboard is maintained and constantly updated [5]. When major players release a new FM (or FM version), they typically advertise their score on SWE-V or one of its variations. That is, SWE-V has become the de facto SE benchmark. In terms of role (ii), researchers have leveraged the rich code and problem solving logic in SWE-V to train and improve models as well as create autonomous agents. A prominent example is SkyRL [6], which is an open-source reinforcement learning framework developed by the Berkeley Sky Computing Lab to train AI agents for long-horizon tasks in real-world environments. SkyRL leverages SWE-V as both the environmental scaffold and training signal for reinforcement learning. Beyond SkyRL, recent evidence shows that SWE-bench-like datasets can even enhance the reasoning capabilities of general-purpose FMs for everyday tasks [7, 8].
@@
While SWE-bench-like datasets have vast practical value, they are also exceptionally costly and difficult to construct. SWE-V was created by OpenAI as a response to the original SWE-bench, which had incomplete issue reports, issue reports containing the reference solution, and unit tests checking for aspects not even described in the issue. These data problems are critical as they lead to unsolvable issues, since either the problem to be solved cannot be properly understood (not even by a human) or candidate solutions generated by the model cannot be adequately verified (e.g., a valid solution may fail the tests). SWE-V excludes unsolvable tasks from the original dataset, focusing solely on those that are verifiably solvable. This was achieved through manual and careful labeling of issues (are issues clear enough?) and test patches (is test coverage adequate and precise?) of 1,699 randomly drawn instances from SWE-bench. We estimate that constructing SWE-V (500 instances) required approximately 2,265.3 engineer hours of manual labeling. The labeling work was performed by professional software developers hired specifically for the task, at an estimated cost exceeding \$170,000111All reported monetary values in this paper are in US dollars (USD).. Additional details are provided in RQ3 (Section V).
@@
In summary, manual labeling of SWE-bench-like datasets is time-consuming, expensive, subject to labeler bias, and does not scale. Compounding these challenges, the scarcity of high-quality, manually labeled benchmarks has driven the community to rely repeatedly on the same SWE-bench data—raising the risk of overfitting and potentially training on the ben hampering meaningful progress. Given that LLM-makers quite potentially train new LLMs on the benchmarking data [12], fresh benchmarks are required. To address these challenges, we present SPICE[Uncaptioned image]: an automated platform for labeling SWE-bench-like datasets that strives for data quality at scale. Leveraging Aider [13], a context-aware code navigation and editing AI assistant, SPICE extracts relevant information from repositories, constructs structured prompts, and synthesizes high-quality labels for issue clarity and test coverage in a scalable manner. To enhance robustness, SPICE labels each instance multiple times and subsequently performs an aggregation through majority voting (a.k.a., self-consistency via stochastic sampling [14]). This process yields a more statistically grounded approximation of expert consensus.
The development of SPICE was informed by our own experience in manually labeling a SWE-bench-like dataset (SWE-Gym [2]). Over the course of annotating over 800 instances, we encountered persistent subjectivity, low IRA, and significant cognitive and time demands, even with a team of expert annotators (see Section III). These challenges made clear that manual labeling at scale is neither sustainable nor reproducible. SPICE emerges from this hard-earned experience as a principled response. As a demonstration of SPICE’s scalability, we release SPICE[Uncaptioned image]-bench—a curated dataset of 6,802 automatically labeled instances drawn from 291 real-world open-source projects in SWE-Gym [2]. This is the largest known collection of solvable SWE-bench-like tasks, and is over 13 times larger than SWE-V’s human-labeled subset. SPICE-bench provides a rich resource for fine-tuning, benchmarking, and training SE-focused foundation models.
@@
  • RQ1: How accurately does SPICE reproduce expert-curated SWE-V labels?
    Spice achieves an accuracy of 67% in identifying “unsolvable” instances from the SWE-bench dataset.
  • RQ2: How similar are the SPICE-generated rationales to those in SWE-V?
    The rationales produced by SPICE’s labelers are semantically similar to that of expert labelers and are sufficiently detailed to be debuggable.
@@
    Spice reduces the labeling cost of 1,000 instances from an estimated \$100,000 (using expert software engineer labelers) to only \$5.10.
@@
  • We release a dataset of 6,802 SPICE labeled SWE bench instances, which is the largest curated collection of solvable SWE-bench-like instances till date. Our dataset is available on Hugging Face as SPICE-bench*222*The URL will be made available after acceptance, but the dataset and our SPICE[Uncaptioned image]tool is enclosed with the submission as supplementary material.
@@ II SWE-Bench Verified (SWE-V)
SWE-V was introduced in mid-2024 as a solution to SWE-bench’s validity issues [1]. This subset of 500 instances was selected through an intensive labeling campaign aimed at removing the flawed instances (unsolvable issues)  [18]. The campaign recruited a total of 93 professional Python developers, who labeled 1,699 random instances from SWE-bench. Every instance was labeled by 3 developers. More specifically, labelers were assigned two fundamental assessment tasks: Issue Clarity Assessment (ICA) and Test Coverage Assessment (TCA). An instance was deemed verified when it passed both assessments. Additionally, as supplementary information (i.e., not used for dataset filtering), labelers were asked to estimate how long it would take an experienced software engineer, after spending a few hours getting familiar with the codebase, to write a patch that resolves the issue. Since this effort estimation does not impact the filtering, we do not discuss it in this paper (however, SPICE does provide an estimation for effort as well). In the following, we describe ICA and TCA. More details can found in SWE-V webpage [1].
@@ III From Manual Effort to Automation: Lessons from Replicating SWE-bench-Verified
SWE-V has emerged as the de facto standard for evaluating FMs on realistic software maintenance tasks. However, it remains unclear whether its labor-intensive annotation protocol is replicable by independent teams. To tackle this problem, we conducted a manual annotation campaign modeled closely after SWE-V’s rubric, targeting 840 unlabeled instances from the SWE-Gym dataset - an unlabeled, SWE-bench-like benchmark - to curate our own dataset.
@@ III-C Results, Lessons Learned, and Key Takeaways
Key Takeaway. Without involvement from developers who actively contribute to each project, large-scale manual labeling for SWE-V-like datasets is impractical.
@@ IV-A Issue Clarity Assessment (ICA) Pipeline
Rationale-Informed Prompt. Standard prompting strategies such as zero-shot, few-shot, chain-of-thought, and domain informed checklist-based approaches failed to consistently produce high-quality issue clarity labels. Inspired by Lin et al. [25], who showed that human feedback improves prompt effectiveness, we grounded our prompt in developer rationales provided by SWE-V annotators.
We selected OpenAI-labeled issues where all three annotators agreed on the label. From this filtered pool, we performed stratified sampling ($n=5$ per clarity level; see Section II) to capture a diverse set of 20 rationales across issues with different clarity levels. For each selected issue, we extracted the title, description, label, and all three rationales provided by the SWE-V labelers (see Section II). We analyzed these with DeepSeek-R1 [26] using exploratory prompting to identify recurring rationale patterns, such as missing context, vague language, or lack of reproducibility and how they related to an issue being labeled well-specified or underspecified. These patterns helped us create a rationale checklist that guided prompt construction.
@@ V SPICE Evaluation Results
V-A RQ1: How accurately does SPICE reproduce expert-curated SWE-V labels?
@@
Approach. To evaluate how accurately SPICE reproduces expert-curated SWE-V labels, we adopt a two-pronged strategy. First, we perform an automatic evaluation by computing SPICE’s accuracy for both ICA and TCA tasks against SWE-V labels (SWE-V Accuracy Evaluation). Second, we conduct a manual evaluation, where two labelers independently assess whether SPICE labels are correct and whether the provided rationale is appropriate. In cases of disagreement, a third labeler adjudicates to determine the final label. We detail our evaluation setup below.
SWE-V Accuracy Evaluation.
@@
– Evaluation Set. We sampled 110 evaluation instances using quota sampling, selecting 10 instances from each of the 12 projects in the SWE-V dataset [28]. This strategy enables us to assess the generalizability of SPICE across a diverse range of projects on both ICA and TCA tasks. For two projects with fewer than 10 available instances (flask had only one instance and seaborn had nine), we selected all available instances, resulting in a final total of 110. While 110 instances may appear small at first glance, it exceeds the 91 instances that would result from randomly sampling SWE-V with a 95% confidence level and a 10% error margin, a common practice [29]. Given these factors, we consider our 110-instance evaluation set to be both robust and justified for this study.
@@
For these instances, we did not use the final ICASWE-V-label and TCASWE-V-label provided by SWE-V moderators. Instead, we derived reference labels via majority vote across the individual ICA and TCA annotations from SWE-V labelers. This approach mitigates potential moderator bias [9, 10] and ensures a fairer comparison, as SPICE also uses majority voting to determine its ICA and TCA labels.
@@
– FM Selection. To ensure a representative evaluation, we selected a diverse mix of open- and closed-source FMs, including both reasoning-oriented and general-purpose models. We include open-source models like DeepSeek R1 Reasoner [30] and closed-source models from OpenAI [31]: GPT-4o, GPT-4o mini, and GPT-4.1. We also evaluate locally hosted models using the Ollama framework [32], including DeepSeek 32B (Ollama-Deepseek-R1:32b), Qwen-14B (Ollama-Qwen2.5:14b), and Qwen-7B (Ollama-Qwen2.5:7b). This setup spans a range of parameter sizes and deployment contexts (covering both locally hosted and externally hosted environments) allowing us to comprehensively assess SPICE’s performance across varied FM capabilities.
@@
Finally, we ran SPICE on all 110 evaluation instances, computed its ICASPICE-label and TCASPICE-label, and measured their accuracy against the corresponding SWE-V reference labels.
Manual Agreement-based Evaluation.
@@
We conducted a complementary manual evaluation to directly assess the validity of SPICE generated labels using the best performing FM from the previous step and their associated rationales. As noted in Section III, we observed low IRA among labelers in the SWE-V dataset. Therefore, while SPICE may exhibit high accuracy on the SWE-V evaluation, such scores may not fully reflect its real-world usefulness. In this setup, labelers reviewed the ICASPICE-label and TCASPICE-label along with their corresponding rationales, and responded to a simple but structured questionnaire.
@@
SWE-V Accuracy Evaluation Results.
SPICEG-4o-m achieves an accuracy of 87.3% on ICA, while SPICEDS-R achieves 68.5% on TCA. Table I summarizes SPICE’s performance across all studied FMs on both ICA and TCA tasks. As shown, SPICE consistently performs better on ICA than on TCA. Even the lowest-performing model, SPICEQ-7 (which we were unable to evaluate on TCA due to segmentation faults in Aider) achieves an accuracy of 68.2% on ICA, which matches our best TCA score. We hypothesize that this discrepancy arises from the inherent complexity of the TCA task. Accurate TCA labeling requires SPICE to process and reason over significantly more contextual information, as discussed in Section IV, which likely exceeds the capabilities of smaller models.
Manual Agreement-based Evaluation Results.
Our evaluation found that SPICE correctly labeled ICA 93.5% of the time, and TCA 60% of the time. These results suggest that SPICE can be reliably adopted for ICA, but that TCASPICE-label should be used with caution and accompanied by human review.
@@
V-B RQ2: How similar are the SPICE-generated rationales to those in SWE-V?
Motivation. Although Section III shows that SWE-V labels suffer from low IRA, the benchmark remains popular because those labels are accompanied by a human-written rationale. These rationales make post-hoc verification and debugging feasible, which is critical for a dataset that is intrinsically hard to label. For SPICE to gain similar trust and adoption, its FM-generated rationales must approach the quality of those written by SWE-V labelers. We therefore measure how closely SPICE rationales resemble human rationales.
Approach. To answer this RQ, we use the same 110 instances and assess the semantic similarity between the rationales written by Human labelers and SPICE labelers. We treat each of the three runs in SPICE as equivalent to a SWE-V labeler and cover rationales for both ICA and TCA. To enable consistent comparison across annotations, we abstract SWE-V annotators into three rotating “seats” (Human L1-L3), representing the three annotator roles per instance. These seats are filled by different subsets of the 93 human annotators in SWE-V, allowing us to compute semantic similarity without identity mismatches across instances (see Section II).
To compute semantic similarity, we embed the three SWE-V rationales per instance into a single document using the text-embedding-3-large [35] embedding model. We do this both individually, and all three human rationales combined. We apply the same embedding procedure to SPICE rationales. Next, we compute the cosine similarity between the resulting embeddings (humans vs SPICE) for comparison. To ensure that SPICE-produced rationales are neither overly complex nor simplistic, we compare the word counts of SPICE and SWE-V rationales. For this RQ, we use SPICEG-4o-m for ICA and SPICEDS-R for TCA, as they demonstrated the best performance in the previous RQ.
Results. SPICE rationales exhibit strong semantic alignment with human rationales, achieving a median similarity of 0.72 for ICA and 0.743 for TCA (Figure 2, 3). These results suggest that SPICE labelers consistently produce explanations that reflect human reasoning.

### v5

@@ I Introduction
In the era of Foundation Models (FMs) like Large Language Models (LLMs), high-quality software engineering (SE) datasets fulfill two key roles: (i) they benchmark model performance on real-world SE tasks and (ii) serve as valuable corpora for pretraining and fine‑tuning FMs.
@@
A prominent example of a high-quality SE dataset is SWE-bench (SoftWare Engineering benchmark) Verified [1] (SWE-V). SWE-V was designed to assess an FM’s ability to produce a valid code patch that fixes an issue report from a given software repository. More specifically, each instance (row) in SWE-V represents a code repair task and contains three key fields: the issue description (title and body), a reference (golden) solution patch, and corresponding test cases used for evaluation. We refer to similar SE datasets containing these key fields as SWE-bench-like datasets. Notable SWE-bench-like datasets include SWE-Gym [2], Multi-SWE‑bench [3], and SWE-bench Multimodal [4]. The tasks in SWE-V span 12 real-world Python projects, including widely used ones such as Django.
@@
In terms of role (i), SWE-V has been extensively used to benchmark FMs as well as agentic scaffolds built on top of those. A leaderboard is maintained and constantly updated [5]. When major players release a new FM (or FM version), they typically advertise their score on SWE-V or one of its variations. That is, SWE-V has become the de facto SE benchmark. In terms of role (ii), researchers have leveraged the rich code and problem solving logic in SWE-V to train and improve models as well as create autonomous agents. A prominent example is SkyRL [6], which is an open-source reinforcement learning framework developed by the Berkeley Sky Computing Lab to train AI agents for long-horizon tasks in real-world environments. SkyRL leverages SWE‑V as both the environmental scaffold and training signal for reinforcement learning. Beyond SkyRL, recent evidence shows that SWE-bench-like datasets can even enhance the reasoning capabilities of general-purpose FMs for everyday tasks [7, 8].
@@
While SWE-bench-like datasets have vast practical value, they are also exceptionally costly and difficult to construct. SWE-V was created by OpenAI as a response to the original SWE-bench, which had incomplete issue reports, issue reports containing the reference solution, and unit tests checking for aspects not even described in the issue. These data problems are critical as they lead to unsolvable issues, since either the problem to be solved cannot be properly understood (not even by a human) or candidate solutions generated by the model cannot be adequately verified (e.g., a valid solution may fail the tests). SWE-V excludes unsolvable tasks from the original dataset, focusing solely on those that are verifiably solvable. This was achieved through manual and careful labeling of issues (are issues clear enough?) and test patches (is test coverage adequate and precise?) of 1,699 randomly drawn instances from SWE-bench. We refer to this dataset as SWE-L. We estimate that constructing SWE-V (500 instances) required approximately 2,265.3 engineer hours of manual labeling. The labeling work was performed by professional software developers hired specifically for the task.
@@
In summary, manual labeling of SWE-bench-like datasets is time-consuming, expensive, subject to labeler bias, and does not scale. Compounding these challenges, the scarcity of high-quality, manually labeled benchmarks has driven the community to rely repeatedly on SWE-V, considerably raising the risks of overfitting and benchmark data leakage  [12] (i.e., models being evaluated with data that they have seen during pretraining). Consequently, it is crucial to establish a streamlined approach for the rapid and reliable generation of high-quality SWE-bench-like datasets. To address these challenges, we present SPICE[Uncaptioned image]: an automated platform for labeling SWE-bench-like datasets that strives for data quality at scale. Leveraging Aider [13], a context-aware code navigation and editing AI assistant, SPICE extracts relevant information from repositories, constructs structured prompts, and synthesizes high-quality labels for issue clarity and test coverage in a scalable manner. To enhance robustness, SPICE labels each instance multiple times and subsequently performs an aggregation through majority voting (a.k.a., self-consistency via stochastic sampling [14]). This process yields a more statistically grounded approximation of expert consensus.
The development of SPICE was informed by our own experience in manually labeling a SWE-bench-like dataset (SWE‑Gym [2]). Over the course of annotating over 800 instances, we encountered persistent subjectivity, low IRA, and significant cognitive and time demands, even with a team of expert annotators (see Section III). These challenges made clear that manual labeling at scale is neither sustainable nor reproducible. SPICE emerges from this hard-earned experience as a principled response. As a demonstration of SPICE’s scalability, we release SPICE[Uncaptioned image]-bench—a curated dataset of 6,802 automatically labeled instances drawn from 291 real-world open-source projects in SWE-Gym [2]. This is the largest known collection of solvable SWE-bench-like tasks, and is over 13 times larger than SWE-V’s human-labeled subset. SPICE-bench provides a rich resource for fine-tuning, benchmarking, and training SE-focused foundation models.
@@
  • RQ1: How accurately does SPICE reproduce expert-curated SWE-L labels?
    SPICE achieves an accuracy of 67% in identifying the instances that were labeled as “unsolvable” in SWE-L.
  • RQ2: How similar are the SPICE-generated rationales to those in SWE-L?
    The rationales produced by SPICE’s labelers are semantically similar to those of expert labelers and are sufficiently detailed to be debuggable.
@@
    In its default setting, SPICE costs only \$5.10 to label 1,000 instances [^1] All reported monetary values in this paper are in US dollars (USD)..
@@
  • We release a dataset of 6,802 SPICE-labeled, which is the largest curated collection of solvable SWE-bench-like instances till date. Our dataset is available on Hugging Face as SPICE-bench*222*The URL will be made available after acceptance, but the dataset and our SPICE[Uncaptioned image]tool is enclosed with the submission as supplementary material.
@@ II SWE-Bench Verified (SWE-V)
SWE-V was introduced in mid-2024 as a solution to SWE-bench’s validity issues [1]. This subset of 500 instances was selected through an intensive labeling campaign aimed at removing the flawed instances (unsolvable issues)  [18]. The campaign recruited a total of 93 professional Python developers, who labeled 1,699 random instances from SWE-bench. We refer to this dataset as SWE-L (SWE-Bench labeled). Every instance was labeled by 3 developers. More specifically, labelers were assigned two fundamental assessment tasks: Issue Clarity Assessment (ICA) and Test Coverage Assessment (TCA). An instance was deemed verified when it passed both assessments. Additionally, as supplementary information (i.e., not used for dataset filtering), labelers were asked to estimate how long it would take an experienced software engineer, after spending a few hours getting familiar with the codebase, to write a patch that resolves the issue. Since this effort estimation does not impact the filtering, we do not discuss it in this paper (however, SPICE does provide an estimation for effort as well). In the following, we describe ICA and TCA. More details can found in SWE-V webpage [1].
@@ III From Manual Effort to Automation: Lessons from Replicating SWE-bench-Verified
SWE-V has emerged as the de facto standard for evaluating FMs on realistic software maintenance tasks. However, it remains unclear whether its labor-intensive annotation protocol is replicable by independent teams. To tackle this problem, we conducted a manual annotation campaign modeled closely after SWE-V’s rubric, targeting 840 unlabeled instances from the SWE-Gym dataset - an unlabeled, SWE-bench-like dataset - to curate our own dataset.
@@ III-C Results, Lessons Learned, and Key Takeaways
Key Takeaway. Without involvement from developers who actively contribute to each project, large-scale manual labeling for SWE-bench-like datasets is impractical.
@@ IV-A Issue Clarity Assessment (ICA) Pipeline
Rationale-Informed Prompt. Standard prompting strategies such as zero-shot, few-shot, chain-of-thought, and domain informed checklist-based approaches failed to consistently produce high-quality issue clarity labels. Inspired by Lin et al. [25], who showed that human feedback improves prompt effectiveness, we grounded our prompt in developer rationales provided by SWE-L annotators.
We selected OpenAI-labeled issues where all three annotators agreed on the label. From this filtered pool, we performed stratified sampling ($n=5$ per clarity level; see Section II) to capture a diverse set of 20 rationales across issues with different clarity levels. For each selected issue, we extracted the title, description, label, and all three rationales provided by the SWE-L labelers (see Section II). We analyzed these with DeepSeek-R1 [26] using exploratory prompting to identify recurring rationale patterns, such as missing context, vague language, or lack of reproducibility and how they related to an issue being labeled well-specified or underspecified. These patterns helped us create a rationale checklist that guided prompt construction.
@@ V SPICE Evaluation Results
V-A RQ1: How accurately does SPICE reproduce expert-curated SWE-L labels?
@@
Approach. To evaluate how accurately SPICE reproduces expert-curated SWE-L labels, we adopt a two-pronged strategy. First, we perform an automatic evaluation by computing SPICE’s accuracy for both ICA and TCA tasks (SWE-L-Based Accuracy Evaluation). Second, we conduct a manual evaluation, where two labelers independently assess whether SPICE labels are correct and whether the provided rationale is appropriate. In cases of disagreement, a third labeler adjudicates to determine the final label. Finally, to better understand the quality of SPICE-generated labels, we conduct an experiment in which we determine whether there is a relationship between the number of top scaffolds that solved a given SWE-V instance and the ICA and TCA labels produced by SPICE (Scaffold-based Evaluation). We detail our evaluation setup below.
SWE-L-based Accuracy Evaluation:
@@
– Evaluation Set. We sampled 110 evaluation instances using quota sampling, selecting 10 instances from each of the 12 projects in the SWE-L dataset [28]. This strategy enables us to assess the generalizability of SPICE across a diverse range of projects on both ICA and TCA tasks. For two projects with fewer than 10 available instances (flask had only one instance and seaborn had nine), we selected all available instances, resulting in a final total of 110 instances. While 110 instances may appear small at first glance, it exceeds the 91 instances that would result from randomly sampling SWE-L with a 95% confidence level and a 10% error margin (a common practice in Empirical SE  [29]). Given these factors, we consider our 110-instance evaluation set to be both robust and justified for this study.
@@
For these instances, we did not use the final ICASWE-L-label and TCASWE-L-label provided by SWE-L moderators. Instead, we derived reference labels via majority vote across the individual ICA and TCA annotations from SWE-L labelers. This approach mitigates potential moderator bias [9, 10] and ensures a fairer comparison, as SPICE also uses majority voting to determine its ICA and TCA labels.
@@
– FM Selection. To ensure a representative evaluation, we selected a diverse mix of open- and closed-source FMs, including both reasoning-oriented and general-purpose models. We include open-source models like DeepSeek R1 Reasoner [30] and closed-source models from OpenAI [31]: GPT-4o, GPT‑4o mini, and GPT‑4.1. We also evaluate locally hosted models using the Ollama framework [32], including DeepSeek 32B (Ollama-Deepseek-R1:32b), Qwen‑14B (Ollama-Qwen2.5:14b), and Qwen‑7B (Ollama-Qwen2.5:7b). This setup spans a range of parameter sizes and deployment contexts (covering both locally hosted and externally hosted environments) allowing us to comprehensively assess SPICE’s performance across varied FM capabilities.
@@
Finally, we ran SPICE on all 110 evaluation instances, computed its ICASPICE-label and TCASPICE-label, and measured their accuracy against the corresponding SWE-L reference labels. Note that we use SWE-L instead of SWE-V since SWE-V only has positive samples (i.e., instances with well-specified issues and adequate test coverage).
Manual Agreement-based Evaluation:
@@
We conducted a complementary manual evaluation to directly assess the validity of SPICE generated labels using the best performing FM from the previous step and their associated rationales. As noted in Section III, we observed low IRA among labelers in the SWE-L dataset. Therefore, while SPICE may exhibit high accuracy on the SWE-L evaluation, such scores may not fully reflect its real-world usefulness. In this setup, labelers reviewed the ICASPICE-label and TCASPICE-label along with their corresponding rationales, and responded to a simple but structured questionnaire.
@@
Scaffold-based Evaluation: To get further insights into the quality of the labels produced by SPICE, we carry out the following experiment. We first visit the SWE-V leaderboard[5] and identify the top-25 scaffolds (we tick the “Checked” filter). Next, we clone the SWE-Bench experiments repository (https://github.com/swe-bench/experiments) and visit the main/evaluation/verified path. This path contains several subdirectories, each representing the prediction results of a scaffold shown in the leaderboard webpage. After manually mapping the names displayed in the leaderboard to the subfolder names, we determine the list of “resolved” instances for each scaffold. Next, we compute a metric called resolution rate for each SWE-V instance, which we define as the fraction of scaffolds that resolved that instance (x / 25). We now run SPICE for all SWE-V instances to compute their ICA and TCA labels. Our hypothesis is that instances that are well-specified and have adequate test coverage will have a higher resolution rate than those that are poorly-specified and have inadequate test coverage. Hence, we test whether ICA and TCA have an effect on the resolution rate using Anova tests (alpha = 0.05). We used GPT-4.1 to perform this experiment on 250 random SWE-V instances.
Results. (SWE-L-based Accuracy Evaluation) SPICEG-4o-m achieves an accuracy of 87.3% on ICA, while SPICEDS-R achieves 68.5% on TCA. We interpret these scores as lower bounds for correctness, since the cases in which SPICE and SWE-L agree with each other are likely to represent “correct” labels more often than not. Hence, we consider our ICA and TCA accuracies to represent solid lower bounds for label correctness (especially considering the pioneering nature of SPICE).
Table I summarizes SPICE’s performance across all studied FMs on both ICA and TCA tasks. As shown, SPICE consistently performs better on ICA than on TCA. Even the lowest-performing model SPICEQ-7 (which we were unable to evaluate on TCA due to segmentation faults in Aider) achieves an accuracy of 68.2% on ICA, which matches our best TCA score. We hypothesize that this discrepancy arises from the inherent complexity of the TCA task. Accurate TCA labeling requires SPICE to process and reason over significantly more contextual information, as discussed in Section IV, which likely exceeds the capabilities of smaller models.
(Manual Agreement-based Evaluation) SPICE correctly labeled ICA 93.5% of the time, and TCA 60% of the time. These results suggest that SPICE can be reliably adopted for ICA, but that TCASPICE-label should be used with caution and accompanied by human review.
@@
(Scaffold-based Evaluation) There is a statistically significant main effect of TCA labels on the resolution rate. We first analyze the effect of TCA on the resolution rate. Instances with adequate test coverage have a higher resolution rate (n = 198, median = 0.44) compared to those with an inadequate test coverage (n = 52, median = 0.32). This difference is further supported by a one-way ANOVA, which indicates a statistically significant effect of the TCA label on the task resolution rate (F = 6.3, p = 0.01), reinforcing the reliability of SPICE labels. The distributions can be seen in Figure 2.
Figure 2: A comparison between the task resolution rate for SWE-V instances with inadequate (green) and adequate (orange) test coverage.
Next we analyze ICA. The SPICE labels indicate that 241 instances were classified as well-specified, while only 9 were classified as underspecified. Since all instances in SWE-V are positive, SPICE achieves high ICA accuracy by closely mirroring this distribution. Because the ICA labels are highly imbalanced (only 9 underspecified cases), assumptions of ANOVA regarding group size and variance homogeneity are difficult to meet, making inference unreliable. In particular, only three instances are both underspecified and have inadequate test coverage, rendering a two-way ANOVA also difficult to interpret. Notably, the lower TCA accuracy (Table 1) introduces greater diversity across labels, which is beneficial for evaluating task resolution rates on the SWE-V leaderboard.
V-B RQ2: How similar are the SPICE-generated rationales to those in SWE-L?
Motivation. Although Section III shows that SWE-L labels suffer from low IRA, the SWE-V benchmark remains popular because those labels are accompanied by a human-written rationale. These rationales make post-hoc verification and debugging feasible, which is critical for a dataset that is intrinsically hard to label. For SPICE to gain similar trust and adoption, its FM-generated rationales must approach the quality of those written by SWE-L labelers. We therefore measure how closely SPICE rationales resemble human rationales.
Approach. To answer this RQ, we use the same 110 instances and assess the semantic similarity between the rationales written by Human labelers and SPICE labelers. We treat each of the three runs in SPICE as equivalent to a SWE-L labeler and cover rationales for both ICA and TCA. To enable consistent comparison across annotations, we abstract SWE-L annotators into three rotating “seats” (Human L1-L3), representing the three annotator roles per instance. These seats are filled by different subsets of the 93 human annotators in SWE-L, allowing us to compute semantic similarity without identity mismatches across instances (see Section II).
To compute semantic similarity, we embed the three SWE-L rationales per instance into a single document using the text-embedding-3-large [35] embedding model. We do this both individually and with all three human rationales combined. We apply the same embedding procedure to SPICE rationales. Next, we compute the cosine similarity between the resulting embeddings (humans vs SPICE) for comparison. To ensure that SPICE-produced rationales are neither overly complex nor simplistic, we compare the word counts of SPICE and SWE-L rationales. For this RQ, we use SPICEG-4o-m for ICA and SPICEDS-R for TCA, as they demonstrated the best performance in the previous RQ.
Results. SPICE rationales exhibit strong semantic alignment with human rationales, achieving a median similarity of 0.72 for ICA and 0.743 for TCA (Figure 3, 4). These results suggest that SPICE labelers consistently produce explanations that reflect human reasoning.